\begin{thebibliography}{10}

\bibitem{algebrization}
Scott Aaronson and Avi Wigderson.
\newblock Algebrization: a new barrier in complexity theory.
\newblock {\em ACM Trans. Comput. Theory}, 1(1):54, 2009.
\newblock Id/No 2.

\bibitem{baker1975relativizations}
Theodore Baker, John Gill, and Robert Solovay.
\newblock Relativizations of the p=?np question.
\newblock {\em SIAM Journal on computing}, 4(4):431--442, 1975.

\bibitem{Hstad2017AnAD}
Johan H{\aa}stad, Benjamin Rossman, Rocco~A. Servedio, and Li-Yang Tan.
\newblock An average-case depth hierarchy theorem for boolean circuits.
\newblock {\em Journal of the ACM (JACM)}, 64:1--27, 2017.

\bibitem{Homer2011}
Steven Homer and Alan~L. Selman.
\newblock {\em Computability and Complexity Theory, Second Edition}.
\newblock Texts in Computer Science. Springer, 2011.

\bibitem{KleeneAH}
S.~C. Kleene.
\newblock Recursive predicates and quantifiers.
\newblock {\em Trans. Am. Math. Soc.}, 53:41--73, 1943.

\bibitem{ko88}
Ker{-}I Ko.
\newblock Relativized polynomial time hierarchies having exactly {K} levels.
\newblock In {\em Proceedings: Third Annual Structure in Complexity Theory
  Conference, Georgetown University, Washington, D. C., USA, June 14-17, 1988}.
  {IEEE} Computer Society, 1988.

\bibitem{Mostowski1947}
Andrzej Mostowski.
\newblock On definable sets of positive integers.
\newblock {\em Fundam. Math.}, 34:81--112, 1947.

\bibitem{ClassRecTh}
Piergiorgio Odifreddi.
\newblock {\em Classical recursion theory. {The} theory of functions and sets
  of natural numbers}, volume 125 of {\em Stud. Logic Found. Math.}
\newblock Amsterdam etc.: North-Holland, 1989.

\bibitem{post1948degrees}
Emil~L Post.
\newblock Degrees of recursive unsolvability-preliminary report.
\newblock In {\em Bulletin of the American Mathematical Society}, volume~54,
  pages 641--642. AMER MATHEMATICAL SOC 201 CHARLES ST, PROVIDENCE, RI
  02940-2213, 1948.

\bibitem{rogers1967theory}
H.~Rogers.
\newblock {\em Theory of Recursive Functions and Effective Computability}.
\newblock Higher Mathematics Series. McGraw-Hill, 1967.

\bibitem{Rossman2015ComplexityTC}
Benjamin Rossman, Rocco~A. Servedio, and Li-Yang Tan.
\newblock Complexity theory column 89: The polynomial hierarchy, random
  oracles, and boolean circuits.
\newblock {\em SIGACT News}, 46:50--68, 2015.

\bibitem{limitlemma}
J.~R. Shoenfield.
\newblock On degrees of unsolvability.
\newblock {\em Annals of Mathematics}, 69(3):644--653, 1959.

\bibitem{soare1987recursively}
R.I. Soare.
\newblock {\em Recursively Enumerable Sets and Degrees: A Study of Computable
  Functions and Computably Generated Sets}.
\newblock Perspectives in mathematical logic. Springer-Verlag, 1987.

\bibitem{compweber}
Rebecca Weber.
\newblock {\em Computability theory}, volume~62 of {\em Stud. Math. Libr.}
\newblock Providence, RI: American Mathematical Society (AMS), 2012.

\bibitem{wegener2005complexity}
Ingo Wegener.
\newblock {\em Complexity Theory: Exploring the Limits of Efficient
  Algorithms}.
\newblock Springer Science \& Business Media, 2005.

\bibitem{Yao1985SeparatingTP}
Andrew Chi-Chih Yao.
\newblock Separating the polynomial-time hierarchy by oracles.
\newblock {\em 26th Annual Symposium on Foundations of Computer Science (sfcs
  1985)}, pages 1--10, 1985.

\end{thebibliography}

\documentclass[runningheads]{llncs}
\newif\ifaa

\usepackage{mypaper}

\usepackage{enumitem} 
\usepackage{verbatim}
\def\llncs
\usepackage{algorithm}
\usepackage{algorithmic}
\ifdefined\lipics
\fi

\newcommand{\nomore}[1]{\ensuremath{\le #1}}
\newcommand{\defeq}{\ensuremath{ := }}

\title{The Polynomial Hierarchy does not collapse}
\makeatletter%
\ifdefined\lipics
\titlerunning{PH does not collapse}

\ccsdesc[500]{Theory of computation~Complexity classes}
\keywords{Complexity classes, Polynomial Hierarchy, Rice's Theorem, P versus NP}
\else
\ifllncs
\author{Reiner Czerwinski}
\institute{TU Berlin(Alumnus)}
\else
\author{Reiner Czerwinski\\TU Berlin (Alumnus)\\reiner.czerwinski@posteo.de}
\fi
\titlerunning{PH does not collapse}
\fi

\newcommand{\B}{\ensuremath{\{0,1\}}}

\begin{document}
\begin{headdata}
  \begin{abstract}

      The arithmetical hierarchy ($\AH$) is similar to the polynomial hierarchy ($\PH$).
      Unlike the $\PH$, the $\AH$ does not collapse relative to any oracle.
      A language in the $(k+1)$-st level of the $\AH$ is computable enumerable (c.e.) relative to the
      $k$th level. So, given an oracle in the $k$th level of the $\AH$, we could
      use a black-box search to decide whether the input word is in the language.

      With very large padding arguments, i.e.\ the paddings grow faster
      than any relative to the level $k$ of the $\AH$ computable function,
      we would construct a language contained in the $k+1$ level of $\PH$,
      if we use only a finite set of input words.
      From the oracle in $\AH$, we would construct an analogue oracle at the
      $k$th level of $\PH$.

      For the input words of the finite set, a word is in the language of $\AH$,
      if and only if it is in the language of $\PH$.
      And the input word is in the oracle set of $\AH$, if and only if it is in the oracle
      of $\PH$.
      As in the language of $\AH$, we must apply a black-box search in the
      language of $\PH$.
      So, we would also have exponentially many
      oracle queries in the language of $\PH$. The $\PH$ does not collapse.
\ifllncs
\keywords{polynomial hierarchy \and aritmetical hierarchy \and complexity classes \and black-box complexity \and Turing degrees}
\fi
    \end{abstract}
\end{headdata}
\section{Introduction}
The polynomial hierarchy (\PH) is similar to the arithmetical hierarchy (\AH).
A language
in the $\AH$ is defined by additional variables, which are bounded by alternating existential and
universal quantifiers. These variables are connected with a computational relation.
The number of alternating quantifiers denotes the level in the hierarchy. 
In the $\PH$ the runtime and the lengths of the quantified variables are bounded
by a polynomial of the input length.

The $\AH$ does not collapse.\
According to the
jump theorem~\cite[page 53, Th. 2.3.]{soare1987recursively}, the $\AH$ does not collapse relative to any oracle. 
A proof for the $\PH$ must circumvent the relativization barrier.
There are oracles where the $\PH$ does not collapse relative to them~\cite{Yao1985SeparatingTP}. The $\PH$ does not collapse relative to random oracles~\cite{Hstad2017AnAD}.
A good overview can be found in~\cite{Rossman2015ComplexityTC}.
On the other hand, 
 for every $k\in\nat$, there is an oracle 
for which the polynomial hierarchy collapses on exactly the $k$th level relative to it~\cite{ko88}.



In Section~\ref{ahph}, we recall the definition of the $\AH$ and the $\PH$
and some properties of the $\PH$.

If we had an oracle $A$ with $\Sigma_1\le_T A$, and a language c.e.\ in $A$, then
we could analyze the number of oracle queries with black-box complexity.
We prove this in Section~\ref{se:blackbox} when the oracle is in $\Pi_{k}.$

In Section~\ref{sec:AHPH},  we 
obtain a connection from the $\AH$ to the $\PH$
via very large padding arguments.

In Section~\ref{nocollapse}, we combine 
the padding technique
of Section~\ref{sec:AHPH} with the black-box complexity results
of Section~\ref{se:blackbox}.
We conclude that the $\PH$ does not collapse.

In section \ref{se:rela} we argue
\ why the proof method does not relativize.
\nocite{rogers1967theory}
\newcommand{\rlf}{\ensuremath{\tilde{R}_{(L,k)}}}
\newcommand{\defiff}{\ensuremath{\overset{\mathrm{def}}{\iff}}}
\newcommand{\defgl}{\ensuremath{\overset{\mathrm{def}}{=}}}
\section{Notations and Prerequisites}\label{ahph}
\subsection{Languages}
A language is a set of words. A word is a finite sequence of symbols.
Without loss of generality we can choose $\B$ as the set of symbols. So, the words are in $\B^*$.
The complement of a language $L$ is $\overline{L}=\bigl(\B^*\setminus L\bigr)$.

\subsection{Relations}
If $A_1, \dots, A_n$ are sets,
then $R \subseteq A_1 \times \cdots \times A_n$ is an $n$-ary relation.
If the tuple $(a_1, \dots, a_n) \in  A_1 \times \cdots \times A_n $
is in $R$, we denote $R(a_1, \dots, a_n)$.
We denote $\lnot R(a_1, \dots, a_n)$, if the tuple is not in the relation $R$.

A relation is computable if it is decidable with a Turing machine, whether
a tuple is in the relation or not.
An $n$-ary relation $R$ is c.e., if there exists a sequence $(R_t)$ of $n$-ary computable relations with
\[R(x_1, \dots , x_n) \iff \exists t_0\;\forall t\ge t_0 : R_t(x_1, \dots , x_n) \]
and
\[\lnot R(x_1, \dots , x_n) \iff \forall t : \lnot R_t(x_1, \dots , x_n) .\]

A relation $R$ is co-c.e.\ if $\lnot R$ is c.e.

Computable relations are more difficult to handle than c.e.\ relations.
A relation is c.e., if there exists a Turing machine which accepts the coding
of the tuple if and only if the tuple is in the relation. Finding a computable
TM is much harder.

\newcommand{\inb}[1][\ast]{\ensuremath{\in\B^{#1}}}
\subsection{Arithmetical Hierarchy}
A language $L$ is in $\Sigma_k$, if there exists a computable relation
$R_L$ with
\begin{equation}\label{defAH}
  \begin{split}x \in L \iff
  \exists y_1\in\B^*\; \forall y_2\in\B^* \dots Q_k y_k\in\B^* :\\
  R_L(x,y_1,y_2,\dots,y_k)
\end{split}
\end{equation}
\[
\text{where }Q_i =\begin{cases} \exists \text{ if }i\text{ is odd} \\ \forall \text{ if }i\text{ is even.} \end{cases}
\]

  The class $\Pi_k$ is defined similarly to $\Sigma_k$, except, that
  the first of the alternating quantifiers is universal and not existential.
  A language $L$ is in $\Pi_k$, if $\overline{L}\in\Sigma_k$.

  The Arithmetical Hierarchy is defined as
  \[
\AH = \bigcup_{k\in\nat} \Sigma_k =\bigcup_{k\in\nat} \Pi_k .   
\]

\begin{remark}
In the literature, the variables in the definition are often natural numbers.
This definition of $\AH$ is equivalent, since
there is a bijective computable function from $\nat$ to $\B^*$:
\begin{equation} \label{eq:bin}
\rho(n) = b_{k-1}b_{k-2}\dots b_0
\text{ where } n+1 = \sum_{j=0}^k b_j*2^j\text{ and }b_k=1.
\end{equation}
\end{remark}

\begin{lemma}\label{ce-rel}
  Let the language $L$ be defined with a c.e.\ or co-c.e.\ relation $R$, such that
  \[ L =\{ x \mid Q_1y_1\inb \;Q_2y_2\inb  \dots Q_{k}y_k\inb \; R(x,y_1,y_2,\dots,y_k) \}, \]
  where $Q_1,Q_2,\dots,Q_k$ are alternating existential and universal
  quantifiers with
  \[Q_K =
    \begin{cases}
      \exists & \text{ if }R\text{ is c.e.}\\
      \forall &\text{ if }R\text{ is co-c.e.}\\
    \end{cases}
  \]
  
  If 
  $Q_1$ is existential (universal), then
  $L$ is in $\Sigma_k$($\Pi_k$).
\end{lemma}

\begin{proof}
  \begin{enumerate}[label=(\roman*)]
    \item
  If $R$ is a c.e.\ relation and $Q_k$ is existential, then there exists a sequence of computable
  relations $R_t$
  with \[R(x,y_1,\dots, y_k) \iff \exists t : R_t(x,y_1,\dots, y_k). \]
  So, \begin{eqnarray*}L &=&\{x \mid Q_1 y_1\inb\;
  Q_2 y_2\inb\;\dots \exists y_k\inb\;\exists t\in \nat :
  \\ && R_t(x, y_1, y_2, \dots, y_k)\} \\
  &=&\{x \mid Q_1 y_1\inb\;
Q_2 y_2\in\B^*\;\dots \exists (y_k,\rho(t))\in (\B^*)^2 :\\
                           &&R_t(x, y_1, y_2, \dots, y_k)\}.
      \end{eqnarray*}
      If the quantifier $Q_1$ is existential, then
      $L$ fulfils the property in Equation~(\ref{defAH}).
      In this case $L\in \Sigma_k$.
      If the quantifier $Q_1$ is universal, then $L_R\in \Pi_k$.
    \item
      If $R$ is co-c.e.\ and $Q_k$ is universal, then
  \[ \overline{L} =\{ x \mid \overline{Q}_1y_1\inb\;\overline{Q}_2y_2\inb \dots \overline{Q}_{k}y_k\inb\;\lnot R(x,y_1,y_2,\dots,y_k) \} \]
  with
  \[ \overline{Q} \defeq \begin{cases} \forall & \text{ if }Q=\exists\\
    \exists & \text{ if }Q=\forall\text{.}
  \end{cases}\]
 As proven in (i), $\overline{L}$ is in $\Pi_k$($\Sigma_k$),
which implies, that $L$ is in $\Sigma_k$($\Pi_k$).
    \end{enumerate}
  \end{proof}



  If $L\in\Sigma_k$ with a computable relation $R_L$, that satisfies property
  (\ref{defAH}), we define the relation for $k \ge 1$:
  \begin{equation} \label{defrlf}
   \rlf (x,y) \defiff \forall y_2\B^* \dots Q_k y_k\in\B^* : R_L(x,y,y_2,\dots,y_k)\text{.}
  \end{equation}
  The set $\{ (x, y) \in (\B^*)^2 \mid \rlf(x,y)\}$ is in $\Pi_{k-1}$.
  So the set is computable in $\Sigma_{k-1}$.
  \begin{theorem}\label{limit}
    Let $k$ be a positive integer, and let $L$ be a language in $\Sigma_k$.
    If $\rlf$ is the binary relation in $\Pi_{k-1}$
    with $L=\{ x \mid \exists y\in\B^* \rlf(x,y)  \}$
    as
    defined in Equation \ref{defrlf},
    and $X$ is a finite set of words,
    then there exists an $M$
    with \[
     \text{ for all }x\in X  : x\in L \iff \exists y\in\B^M \rlf (x,y)\text{.}
    \]
    We denote $M=\eta(X,k,L)$.
  \end{theorem}
  \begin{proof}
    Let $X = \{x_1, \dots ,x_n\}$ be a finite set of words. 
    If $L$ is in  $\Sigma_k$, and $\rlf$ is a relation in $\Pi_{k-1}$
    that satisfies the property in Equation \ref{defrlf},
    then for every $x \in X$ there exists an $m_x\in\nat$ with
    \[m \ge m_x \Rightarrow [ x\in L \iff \exists y\in\B^{\le m} \rlf(x,y) ]\text{.} \]
    Now we define
    \[M = \eta(X,k,L) \defeq \max\bigl(\{ m_x \mid x\in X \land \ x\in L    \} \cup \{0\} \bigr) \text{.}\]
    So,
    \[
     \text{ for all }x\in X  : x\in L \iff \exists y\inb[\le M] \rlf (x,y)\text{.}
    \]
  \end{proof}
  \subsubsection{Properties of $\AH$}
  According to Kleene~\cite{KleeneAH} and Mostowski~\cite{Mostowski1947} we know that:
  \begin{itemize}
  \item $L\in \Sigma_k \iff \overline{L}\in \Pi_k$\\
  \item $\Sigma_k \setminus \Pi_k \not=\emptyset$ and $\Pi_k \setminus \Sigma_k \not=\emptyset$. 
  \end{itemize}
  A language is in $\Sigma_{k+1}$, iff it is c.e.\ in $\Sigma_k$ and $\Pi_k$
  by Post's Theorem~\cite{post1948degrees}.
  A language $A$ is $\Sigma_K$-complete, if $B\le_1 A$ for any language $B\in\Sigma_k$. The $k$-th Turing jump $\emptyset^{(k)}$ is $\Sigma_K$-complete.
  According to the Jump theorem~\cite[page 53, Th. 2.3.]{soare1987recursively}, the $\AH$ does not collapse relative to any oracle.
  
\subsection{Polynomial Hierarchy}
\begin{definition}
  If $A$ is a set and $n>0$, then
  \[ A^{\nomore{n}} \defgl \bigcup_{k\le n} A^k\text{.} \]
\end{definition}
The length of a word $x\in A^*$ is $|x|=\min\{n \mid x\in A^{\nomore{n}}\}$.
The word $1^n$ is the unary coding of $n$.


The polynomial hierarchy is $\PH = \bigcup_{k\ge 0} \Sigma_k^p =\bigcup_{k\ge 0}  \Pi_k^p$.

A language $L$ is in $\Sigma_k^p$ or $\Pi_k^p$ if there exists a relation $R_L$ computable in polynomial time, and polynomial $p$, such that
\[\begin{split} x \in L \Leftrightarrow Q_1y_1\in\B^{\le p(|x|)}\; \dots Q_{k}y_{k}\in\B^{\le p(|x|)}\;:\\ 
  R_L(x, y_1, \dots, y_k)\text{.}
\end{split}
\]
There are alternating existential and universal quantifiers,
\[\text{i.e.\ }
Q_i =\begin{cases} \exists \text{ if }i\text{ is odd} \\ \forall \text{ if }i\text{ is even} \end{cases}\text{ for }L\in\Sigma^p_k \text{ and}\]
\[Q_i =\begin{cases} \forall \text{ if }i\text{ is odd} \\ \exists \text{ if }i\text{ is even} \end{cases}\text{ for }L\in\Pi^p_k\text{.}\]

If $M$ is a TM, then the language of $M$ is $L(M)=\{x \mid M\text{ accepts }x\}$.
A set is c.e.\ if it is the language of a TM\@.

The following theorem is an exercise in some books, e.g. Homework 7.11 in~\cite[p. 166]{Homer2011}. 
\begin{theorem}\label{pattern}
  The set
  \[\begin{split}
      S_k = \{ \langle R , x, 1^{m}, 1^t\rangle \mid \\
      \exists y_1\in\B^{\nomore{m}} \forall y_2\B^{\nomore{m}} \dots Q_k y_k\in\B^{\nomore{m}} :\\
      R( x,y_1,y_2,\dots,y_k) \text{ and } R \text{ is computable within }
       t \text{ steps }\} 
    \end{split}
  \]
  is $\Sigma^p_k$-complete.
  \end{theorem}
\begin{proof}
Obviously, $S_k$ is in $\Sigma_k^p$.
If the language $L$ is in $\Sigma_k^p$, then there exist polynomials $p$, $q$, and an $(n+1)$-ary relation $R$ with
\[\begin{split}
    L = \bigl\{ x \mid \exists y_1\in\B^{\le p(|x|)} \forall y_2\B^{\le p(|x|)} \dots Q_k y_k\in \B^{\le p(|x|)} :\\R(x,y_1,y_2,\dots,y_k) \text{ and }R\text{ is computable within }q(|x|)\text{ steps }\bigr\}.
  \end{split}
\]
So, for all $x$, $\langle R,x,1^{p(|x|)},1^{q(|x|)} \rangle \in S_k$.
This implies $L \le^p_m S_k$.
\end{proof}
\section{Black-Box Search and the $\AH$}\label{se:blackbox}

  \subsection{Oracles and computable Permutations}\label{se:group}

  A computable permutation is a computable bijective function from $\nat$ to \nat,
  or from $\B^*$ to $\B^*$.
The set of computable permutations is a group~\cite[page 73, Th. I]{rogers1967theory}.
So, the inverse of a computable permutation is also a computable permutation.

If $A\subseteq \B^*$ and $f$ is a computable permutation of $\B^*$, then
\[A_f \defeq \{ x \in \B^* \mid f(x)\in A \}. \]
It is obvious, that $A_f \equiv_T A$.
\begin{lemma}\label{swap}
  Let $A$ be a language.
  For every $x,y\in\B^*$ there exists a computable permutation $f$
  with $y=f(x)$.
  Thus
  \[
   x\in A \iff y\in A_f.
 \]
\end{lemma}
\begin{proof}
   The permutation, that swaps $x$ and $y$, i.e.
   \[
     f(w)=
     \begin{cases}
       y & \text{if } w=x \\
       x & \text{if } w=y\\
       w & \text{otherwise,}
     \end{cases}
   \]
   is computable.
 \end{proof}
 
 \subsection{Black-Box Complexity with Oracles}\label{se:bbo}
 \newcommand{\orl}{\ensuremath{\tilde{R}_{L}}}
 \newcommand{\y}{\ensuremath{\overline{y}}}
 \begin{theorem}\label{blackbox}
   Let $k$ and $m$ be positive integers, $ k \ge 1$.
   Let $\tilde{R_L}$ be a $\Pi_{k}$-complete $2$-ary relation. An oracle
   Turing mashine (OTM)
   with an oracle for $\orl$, that decides whether
   $x\in L_m=\{ x \mid \exists y\in\B^{\le m} \orl(x,y)\}$
   needs an oracle query for every $y\in\B^{\le m}$ in the worst case.
 \end{theorem}

 \begin{proof}
   The relation $\orl$ is not computable.
   We assume that we would have an \orl-oracle to search for
   a $y$ with $\orl(x,y)$.
   In this case, we should use black-box search.
   If there is a $\hat{y}\in\B^{\le m}$ with $\orl(x,\hat{y})$, then
   by Lemma~\ref{swap}
   we can construct for each $\y\in\B^{\le m}$ a computable permutation
   $f_{\y}$ with $f_{\y}(\y)=\hat{y}$. The language
   \[ L_{m,f_{\y}} =\{ x \mid \exists y\inb[\le m]\; (\orl\circ f_{\y})(x,y)  \} := \{ x \mid \exists y\inb[\le m]\; \orl(x,f_{\y}(y)) \}\]
   is $(\orl\circ f_{\y} )$-computable, because $(\orl\circ f_{\y}) \equiv_T \orl$. The search result is now at $(x,\y)$.
   So, we have to use the oracle $\tilde{R_L}$ on any
   $y$ with a length less than $m$ in the worst case.
 \end{proof}

 We need a black-box search to decide with the oracle $\orl$
 whether an input word is in
 $L_m =\{ x \mid \exists y\in\B^{\le m} \; \orl(x,y) \}$.

The cardinality of $\bigl(\B^{\le m}\bigr)$ is $N=2^{m+1} - 1$.
If there is only one $y\in\B^{\le m}$ with $\tilde{R}_L(x,y)$, then the black-box complexity,
i.e. the expected number of oracle queries,
is $N/2 +1/2 = 2^m $~\cite[Th. 9.3.1]{wegener2005complexity}.

\section{The Relationship between $\PH$ and $\AH$}\label{sec:AHPH}
\newcommand{\qsizeone}[1][X]{\ensuremath{\mu(#1;k,R_L)}}
\newcommand{\qsizeonepp}[1][X]{\ensuremath{\mu(#1;k+1,R_L)}}
A set of $\AH$ can be transformed into a problem of $\PH$ by considering only a finite set of input words.
For a language $L$ in $\Sigma_k$ we use very large padding arguments to
construct a language $L'$ in $\Sigma_k^p$. For the finite set of input,s
a word is in $L$ ,if and only if the word is in $L'$.
\begin{lemma}\label{finite}
  Let $k$ be a non-negative integer.
  Let $L$ be a language in $\AH$ with c.e.\ or co-c.e.\
  $k$-ary
  relation $R_L$,
  such that
  \[  L = \{ x \mid  Q_1y_1\in\B^{*}\dots Q_{k}y_k\in\B^{*}\;
    R_L(x ,y_1 ,\dots ,y_k) \}
\]
where $Q_1,\dots, Q_k$ are existential or universal quantifiers.
If the set of words $S$ is finite, then
there exist
positive integers $M$ and $T$ with
\[ x\in S \Rightarrow
  [ x \in L \iff  Q_1y_1\in\B^{\nomore{M}}\dots Q_{k}y_k\in\B^{\nomore{M}}
    R_L(x ,y_1 ,\dots ,y_k)
  ]
\]
and $R_L$ is computable within $T$ steps on $\bigl(S \times (    \B^{\le M}  )^k\bigr)$.

We denote $m=\qsizeone$.
\end{lemma}
  \begin{proof}
    We prove this lemma by induction over $k$:
    
      First, we prove the base case $k=0$.
      The relation
      $R_L$ is c.e.\ or co-c.e.
      There is a series of computable unary relations $R_{L,t}$
      with $\lim_t R_{L,t} =R_L $ by the Limit Lemma~\cite{limitlemma}.
      W.l.o.g.\ we can assume, that $R_{L,t}$ is computable within $t$ steps.
      The set $S$
      is finite. Thus,
      for some $T\in\nat$ is $R_{L,T}(x) \iff R_L(x)$ for $x\in S$.
      
      In the induction step, we define the language $L$ with $(k+1)$
      quantifiers and a c.e.\ or co-c.e.\ $(k+2)$-ary relation:
      \[\begin{split}
        L := \{ x \mid Q_1 y_1 \in \B^*
        Q_2 y_2 \in \B^* \dots Q_{k+1}y_{k+1} \in \B^* \\
        R_L(x,y_1,y_2,\dots y_{k+1}) \}
      \end{split}
    \]
    W.l.o.g.\ the quantifier $Q_1$ is existential. Otherwise, we analyze
    $\overline{L}$.

    For convenience, we define the set
    \[
      A := \{ (x,y) \mid Q_2 y_2 \in \B^* \dots Q_{k+1}y_{k+1}\in \B^*\;
      R_L(x,y,y_2,\dots,y_{k+1}) \} \text{.}
    \]
    Obviously,
    \[ x\in L \iff \exists y\in\B^*\; (x,y)\in A \text{.}  \]
    For a finite set of words $S$, there exists an $m_S$ with
    \[
      x\in S \Rightarrow [ x\in L \iff \exists y\in\B^{\le m_S} \text{.}  ]
    \]

    The set $\bigl( S \times \B^{\le m_S} \bigr)$ is finite, and
    $A$ is in $\Pi_k$.
    By the induction hypothesis we get
    some positive integers $m_A$ and $t$ with 
    \[ (x,y)\in A \iff Q_2y_2\in\B^{\le m_A}\dots Q_{k+1}y_{k+1}\in\B^{\le m_A} R_L(x,y,y_2,\dots,y_{k+1}) \]
    and $R_L$ computable in at most $t$ steps
    for $(x,y)\in\bigl( S \times \B^{\le m_S} \bigr)$.

    If we choose
    $M := \qsizeonepp = \max(m_S,m_A)$, we get
    for $x\in S$:
    \[\begin{split}
      \bigl[
      x\in L \iff
       Q_1 y_1 \in \B^{\le M}
        Q_2 y_2 \in \B^{\le M} \dots Q_{k+1}y_{k+1} \in \B^{\le M}\\
        R_L(x,y_1,y_2,\dots y_{k+1})\text{.}
        \bigr]
      \end{split}
    \]
    The set $ S \times (\B^{\le M})^k$ is finite. Thus, there is an integer $T$
    with \[
      \begin{split}(x,y_1,y_2,\dots,y_{k+1}) \in S \times (\B^{\le M})^k
        \Longrightarrow \\
       \bigl [ R_L(x,y_1,y_2,\dots,y_{k+1}) \iff R_{L,T}(x,y_1,y_2,\dots,y_{k+1})\bigr]
      \end{split}
    \] with $\lim_t R_{L,t} = R_L$ and $R_{L,t}$ is computable in $t$ steps.
  \end{proof}
  \nocite{ClassRecTh}
  The arithmetical hierarchy does not collapse~\cite{KleeneAH},~\cite{Mostowski1947}. According to Post's theorem~\cite{post1948degrees},
  $\Sigma_{k+1}$-complete problems have a higher Turing degree than
  $\Sigma_k$-complete problems, i.e.
  if $A$ is $\Sigma_{k+1}$-complete and $B$ is $\Sigma_{k}$-complete,
  then $A\not\le_T B$.
  This implies, that the length $\qsizeone$ of the quantified variables in Lemma~\ref{finite} can grow faster than any $\Sigma_{k-1}$-computable function.

 \section{The $\PH$ does not collapse}\label{nocollapse}

 To show that the $\PH$ does not collapse,
 we prove
 that for a language in $\Sigma_{k+1}^P$ the runtime
 is exponential relative to $\Sigma_{k}^P$.
 \begin{lemma}\label{main}
   Let $k$ be a positive integer.
            If one computes a language in $\Sigma^p_{k+1}$ with a $\Pi^p_{k}$-oracle,
            then in the worst case one needs exponentially many oracle  queries.
          \end{lemma}
          \begin{proof}
            Let \[
         L= \{ x \mid \exists y_1\in\B^*\; \forall y_2\in\B^* \dots Q_{k+1}y_{k+1}\in\B^* \; R_L(x,y_1,y_2,\dots,y_{k+1}) \}
       \] be a $\Sigma_{k+1}$-complete language, such that the relation
       \[ \orl = \{ (x,y) \mid \forall y_2\in\B^* \dots Q_{k+1}y_{k+1}\in\B^* \; R_L(x,y,y_2,\dots,y_{k+1}) \}
         \}
       \] is $\Pi_k$-complete.

       By Lemma \ref{ce-rel}, the language $L$ is in $\Sigma_{k+1}$, if
        the last quantifier $Q_{K+1}$ is existential (universal)
       and the relation $R_L$
       is c.e. (co-c.e.).
       Now, there is $L = \{ x \mid \exists y\in\B^* \orl(x,y)\} $.
     We define \[L_m \defeq \{ x \mid \exists y\in\B^{\le m} \orl(x,y)\} .\]
   Theorem \ref{blackbox} implies that one needs a black-box search
   to decide with oracle $\orl$, whether $x\in L_m$.
   Thus, one needs $\theta(2^m)$ queries to the oracle $\orl$.

   \newcommand{\rlt}[1][t]{\ensuremath{R_{(L,#1)}}}
   If the relation $R_L$ is c.e.\ or co-c.e., then there exists
   a sequence of computable relations $\rlt$ with
   \[ \lim_t \rlt = R_L\] by the limit lemma~\cite{limitlemma}.

   \newcommand{\AAP}{\ensuremath{A_P}}
   \newcommand{\oAAP}{\ensuremath{\tilde{A}_P}}

   Similar to Theorem~\ref{pattern}, we can construct a set
     \begin{equation} 
       \begin{split}
    \AAP \defeq \{ \langle x, 1^m, 1^t  \rangle \mid
    \exists y_1\in\B^{\le m}
    \forall y_1\in\B^{\le m} \dots Q_{k+1} y_{k+1}\in\B^{\le m}\\
    \rlt(x,y_1,y_2,\dots,y_{k+1})
    \}
  \end{split}
\end{equation}

   in $\Sigma_{k+1}^P$. The related $\Pi_{k}^P$-oracle is
  \[\begin{split}
    \oAAP \defeq \{ \langle x, y, 1^m, 1^t  \rangle \mid
    \forall y_2\in\B^{\le m} \dots Q_{k+1} y_{k+1}\in\B^{\le m}\\
    \rlt(x,y,y_2,\dots,y_{k+1})
    \}.
  \end{split}
\]

   If the set $S$ of input words is finite, then there exists an $ M$ 
   and $T$ 
   with
   \begin{equation}\label{big_MT}
     x\in S \Rightarrow [
         x \in L \iff \langle x, 1^M , 1^T \rangle \in \AAP
      ]
    \end{equation}
    
   by Lemma \ref{finite}.
   Obviously,
   \[ \orl(x,y) \iff \langle x, y, 1^M , 1^T \rangle \in \oAAP \]
   for $x\in S$ and $y\in\B^{\le M}$.

   One would like to test whether an $x\in S$ is in
   $L_M =\{x \mid \exists y\in\B^{\le M}\;\orl(x,y)\}$ with an oracle.
   There is not less runtime with oracle $\oAAP$ than with oracle $\orl$,
   because
   $\orl(x,y) \iff \langle x, y, 1^M , 1^T \rangle \in \oAAP $
   for $(x,y) \in S \times \B^{\le M}$.

   So, we need a black-box search to test, whether $x\in S$ is in $L_M$ with the
   oracle $\oAAP$. By Theorem~\ref{pattern}
   the oracle $\oAAP$ is in $\Pi_{k}^P$.

 \newcommand{\ora}{\ensuremath{\mathcal{R}}}
  
 With the oracle $\orl$, we need a black-box search.
 Now, we have to answer the question
 whether there is an oracle that is faster.
 We can construct such an oracle using the linear speed-up theorem,
 which tests many input words at once.
 But the possible speed-up is limited by a computable function,
 and the parameter $M$ in Equation~\ref{big_MT} can grow faster
 than any computable function.
 So, there are also exponentially many queries with the faster oracle.
   
  Perhaps the numbers $M$ and $T$ are not great enough to fulfil
  the property in
  Equation~\ref{big_MT}. If we could test faster with oracle $\oAAP$
  in this case,
  whether an input is
  in $L_M$,
  then we would have a criterion to decide the property of
  Equation~\ref{big_MT}.
  But this is neither computable nor $\Pi_{k}$-computable.
  Thus, we need also black-box search for small bounds $M$ and $T$.
\end{proof}

          According to Lemma~\ref{main}, we need exponential runtime to compute
          a $\Sigma_{k+1}^P$-complete language with a $\Pi_{k}^P$-oracle.
          So, $\Sigma_{k+1}^P \not\le_T^P \Pi_{k}^P$ for every integer $k>0$. 
          
          \section{Notes on the Relativization Barrier}\label{se:rela}
          The polynomial hierarchy does not collapse. This implies
          $\P\not=\NP$. The proof has to circumvent the
          relativization barrier~\cite{baker1975relativizations}.
          In this paper, we use polynomial-long oracle queries.
          Aaronson and Wigderson gave an example~\cite{algebrization},
          where this proof technique fails to relativize.
          So, this proof does not relativize.


\bibliography{lit}{}
\bibliographystyle{plain}
\appendix
\section{Computability Theory in a Nutshell}
A language is a subset of $\omega$, where $\omega$ is an infinitely countable set
such as $\nat$ or $\B^*$.

A language $A$ is $B$-computable, if an OTM with an oracle tape for the set
$B$ could decide, whether an input is in $A$ or not.
The language $A$ is c.e.\ in $B$, when there is an OTM with oracle
$B$, that accepts all words in $A$. Despite to $B$-computable sets, maybe the OTM
does not terminate for words not in $A$.

For languages $A, B$ we denote $A\le_T B$, if $A$ is $B$-computable.
If there exists a computable injective function $f$ with $ x\in A \Rightarrow f(x) \in B$,
then $A\le_1 B$.
For $r\in\{1,T\}$ we denote $A\equiv_r B$, if $A\le_r B$ and $B\le_r A$.
If $A\le_r B$, but $B\not\le_r A$, then $A<_r B$.
Obviously, $A\le_1 B \Rightarrow A\le_T B$.

The Turing jump of a language $A$ is
\[ A' \defeq \{ e | e \in W_e^A \} \]
The set $W_e^A$ contains all input word, that would be
accepted by the OTM with oracle $A$ with the coding $e$. We denote this coding Gödel number.

The $n$-th Turing jump of $A$ is
\[ A^{(n)} = \begin{cases} A & \text{ if }n=0\\
    (A^{(n-1)})' \text{ otherwise}
  \end{cases}
\]
A set $A$ is computable, if $A\equiv_T \emptyset$ So, $A$ is c.e.,if and only if
$A\le_T \emptyset^{(1)}$. A set $A$ is in $\Sigma_n$, if and only if $A\le_1\emptyset^{(n)}$.

The jump theorem~\cite[page 53, Th. 2.3.]{soare1987recursively} is
a collection of some properties of the Turing jump:
\begin{theorem}
  For all $A,B,C \subseteq \omega$
   \begin{enumerate}
   \item $A'$ is c.e.\ in $A$\\
   \item $\A' \not\le_T A$ \\
   \item If $A$ c.e.\ in $B$ and $B\le_T C$ then $A$ is c.e.\ in $C$ \\
   \item $B \le_T A$ if and only if $B'\le_1 A'$\\
   \item If $B\equiv_T A$ then $B'\equiv_1 A'$ \\
   \item A is c.e.\ in $B$, if and only if A is c.e.\ in $\overline{B}$ \\ 
     
   \end{enumerate}
 \end{theorem}

 A god overview of Computability theory is in \cite{compweber}.

\end{document}